# SpiderCat: A Catalog of Compact Binary Millisecond Pulsars

### v1

@@ Abstract
We present SpiderCat, a multi-wavelength catalog of all publicly known compact binary millisecond pulsars (MSPs) in the Galactic field. These systems, colloquially known as “spiders”, consist of neutron stars in tight orbits with low-mass companions, which are gradually ablated by the pulsar wind. SpiderCat includes both primary subclasses – redbacks and black widows – distinguished by companion mass, as well as candidates and peculiar systems such as transitional, huntsman and tidarren MSPs. As of this initial release, SpiderCat contains 109 entries: 32 redbacks, 49 black widows, 23 redback candidates, and 5 black widow candidates.
In this paper, we compile and summarize key parameters for each system, including spin and orbital properties, and multi-wavelength data from radio, optical, X-ray, and $\gamma$-ray observations. An interactive, publicly accessible web interface (https://astro.phys.ntnu.no/SpiderCAT) enables exploration and visualization of the data.
The rapid growth of the number of known spiders, accelerated by the Fermi-LAT survey and its ability to identify MSPs in $\gamma$-rays, has opened the door to population-level studies. Utilizing SpiderCat, we analyze trends in spin period, orbital period, companion mass, emission properties, and spatial distribution. SpiderCat serves as a dynamic, multi-wavelength repository for this unique class of binary pulsars, facilitating new discoveries and constraints on pulsar evolution, particle acceleration, and the neutron star equation of state.
@@ 1 Introduction
Before the Fermi Gamma-ray Space Telescope was launched in 2008, only seven $\gamma$-ray pulsars had been confidently detected (by the Energetic Gamma-Ray Experiment Telescope, EGRET, between 1991 and 2000), all of which were young and slowly rotating (spin period $P>30$ ms; D. J. Thompson, 2008). The Large Area Telescope (LAT) onboard Fermi, with an effective area nearly ten times larger than EGRET’s, has proven to be an efficient pulsar-discovery machine (W. B. Atwood et al., 2009). Based on 12 years of all-sky survey LAT data, Fermi’s Third Pulsar Catalog (3PC) includes 294 $\gamma$-ray pulsars (and 45 candidates; D. A. Smith et al., 2023). Roughly half of these 3PC $\gamma$-ray pulsars are millisecond pulsars (MSPs): “old”, rapidly spinning ($P<30$ ms) neutron stars in binary systems, with relatively low magnetic fields ($B\sim 10^{8}-10^{9}$ G), that have been “recycled” through the accretion of matter (M. A. Alpar et al., 1982). Because MSPs have a higher Galactic scale height than young pulsars (D. R. Lorimer, 2013), they are more easily discovered in LAT surveys far from the Galactic plane, where $\gamma$-ray diffuse emission does not obscure faint point sources.
@@
Many of the new Fermi-LAT pulsar discoveries have turned out to be compact binary MSPs, characterized by short orbital periods (mostly $P_{\rm b}\lesssim 1$ d) and non-degenerate or semi-degenerate companion stars (see definition in Sec. 1.2). In such compact systems, interactions between the pulsar and its companion can be extreme; the outer layers of the companion star may be ablated leading to occultations or “eclipses” of the radio pulsar (A. S. Fruchter et al., 1988). Such destructive effect has inspired cannibalistic spider nicknames for compact binary MSPs, so they are referred to in the literature (and hereinafter) as spider pulsars, spider MSPs, or simply “spiders” (e.g., M. S. E. Roberts, 2011).
@@
Pulsars represent the most numerous Galactic class of $\gamma$-ray sources. They are characterized by low $\gamma$-ray variability and high spectral curvature, which sets them apart from blazars (the most numerous extragalactic class) that present typically higher variability and power law spectra. These characteristics enable an efficient selection of pulsar candidates among thousands of unidentified Fermi-LAT sources ($>$7000 cataloged sources, with $>$2000 still unidentified; S. Abdollahi et al., 2022a). Thus, the Fermi-LAT point source catalogs have been a treasure trove for spider discoveries over the past fifteen years (Sec. 1.1).
@@
An “elite” subset of compact binary MSPs (belonging to the redback subtype, see below) have been observed in the so-called disk state, where they form an accretion disk and the radio pulsar becomes undetectable (Z. Wang et al., 2009; D. de Martino et al., 2015). These so-called transitional MSPs (tMSPs; labeled ‘tr’ below) show optical disk emission lines and a characteristic X-ray variability pattern known as X-ray mode switching (M. Linares et al., 2014). In the pulsar state, when the radio pulsar is detected, spiders show lower luminosities and orbital modulation (with different origins) throughout the electromagnetic spectrum, from radio to $\gamma$-rays. These properties are used to identify new spiders, as described in Sec. 1.1.
@@
As a consequence of the compact orbits, an intra-binary shock (IBS) is thought to form between the pulsar and companion winds in spider MSPs. This shock produces non-thermal (synchrotron) X-ray emission that is modulated with the orbit (Z. Wadiasingh et al., 2017). The IBS can also be an efficient site for particle acceleration (e.g., A. K. Harding & T. K. Gaisser, 1990), potentially contributing to the cosmic-ray positrons detected on Earth (M. Linares & M. Kachelrieß, 2021). During the Gyr-long recycling process, neutron stars in low-mass X-ray binaries can accrete up to about 1 M (e.g., D. Misra et al., 2025a). As a result, MSPs in general – and spiders in particular – are thought to harbor massive neutron stars. Indeed, several studies have found evidence for massive neutron stars from dynamical measurements of compact binary MSPs (e.g., M. Linares et al., 2018; J. Strader et al., 2019; R. W. Romani et al., 2022).
@@
For the reasons above, the field of compact binary MSPs has experienced a boom in the last decade. We now know of more than 100 spiders, enabling meaningful population studies. While existing pulsar catalogs (Sec. 2.1) focus on radio timing properties, there is no public, up-to-date repository of multi-wavelength information for spider MSPs (to the best of our knowledge). Thus, the time is ripe to collect and analyze the spin, orbit and multi-wavelength properties of all known Galactic field spiders. This is the goal of the present work.
@@ 1.1 Spider Discoveries
Thanks to a flurry of Fermi-driven discoveries, the spider population has grown from a rara avis (with only four systems known before 2008; see Figure 1) into an important subclass of MSPs (109 systems presented here as of 2025). Indeed, spiders now constitute 15–20% of the 552 MSPs currently known in the Galactic disk (E. Ferrara & D. Lorimer, 2025, where the higher end of the range includes candidate systems). The methods that have been used to discover new spiders include:
@@
  • Wide-field radio timing surveys (e.g., the Galactic Plane Pulsar Snapshot survey—GPPS—with the Five-hundred-meter Aperture Spherical Telescope, FAST; P. F. Wang et al. 2025; see also A. J. Faulkner et al. 2004). These surveys typically observe each sky location once or only a few times, with relatively short integrations. Because in spider systems the pulsar is occulted during a large fraction of the orbit (20–60%), these surveys are sub-optimal for discovering them. Follow-up observations are required to measure orbital parameters.
@@
  • $\gamma$-ray-guided radio timing searches (e.g., by the Pulsar Search Consortium using the Green Bank Telescope, GBT; J. W. T. Hessels et al. 2011; P. S. Ray et al. 2012; see also E. D. Barr et al. 2013a). This method targets specific unidentified $\gamma$-ray sources using longer and/or more numerous radio pointings, after selecting pulsar candidates based on their $\gamma$-ray properties (as discussed above).
@@
  • $\gamma$-ray-guided optical searches (e.g., the Compact Binary Pulsar Search, COBIPULSE; M. Turchetta et al. 2024; see also R. W. Romani & M. S. Shaw 2011a; A. K. H. Kong et al. 2012; M. Linares et al. 2017; C. Lu et al. 2025). Compared to the more common binary MSPs with white dwarf companions, spider MSPs with main-sequence and brown dwarf companions to spider MSPs are typically more luminous in the optical-IR band. Their compact orbits result in large-amplitude orbital modulations due to the gravitational distortion and irradiation of the companion star, as viewed at different orbital phases. These features enable optical-IR searches for spider candidates, especially the optically brighter redback subclass (Secs. 1.2 and 2.1), by targeting their characteristic flux (and radial velocity) variations. The candidate sources are typically selected based on $\gamma$-ray properties (low variability and highly curved spectra).
@@
  • Blind $\gamma$-ray searches (e.g., H. J. Pletsch et al. 2012; see also L. Nieder et al. 2020). The Fermi-LAT has been surveying the full sky since 2008, providing a rich dataset for blind pulsar searches. These searches fold $\gamma$-ray photons over a range of spin and orbital parameters to detect pulsations from a given source or sky location. Although computationally intensive, especially for MSPs with unknown orbital parameters, they can be aided by prior knowledge of such parameters, e.g., the orbital period (C. J. Clark et al., 2020). In addition, some spiders have shown orbital modulation (H. An et al., 2020) and eclipses (C. J. Clark et al., 2023a) in their $\gamma$-ray light curves.
@@
  • X-ray searches (e.g., S. Bogdanov & J. P. Halpern 2015; see also F. Coti Zelati et al. 2019). X-ray properties can also help identify new spider systems. In particular, X-ray mode switching has been used to discover new tMSP candidates in the disk state. In the pulsar state, features such as the X-ray luminosity ($\sim 10^{30}-10^{33}$ erg s-1), spectral shape (photon index $\simeq 1.5-3$), or orbital modulation from the IBS (often double-peaked with a flux maximum near pulsar inferior conjunction) can aid in identifying spider systems.
@@ 1.2 Spider Definition and Taxonomy
  • A pulsar spin period $P<30$ ms,
  • An orbital period $P_{\rm b}\lesssim 10$ d (most systems have $P_{\rm b}\lesssim 1$ d, indicative of compact orbits, with very few exceptions).
  • A non-degenerate or semi-degenerate companion star, inferred from: i) the presence of radio eclipses, ii) the detection of X-ray mode switching (in tMSPs), iii) the optical-IR spectra of the companion or accretion disk, and iv) the detection and properties of X-ray/optical orbital modulation (see details above).
Spiders are divided into two broad sub-classes, based on the minimum companion mass ($M_{\rm c,min}$):
@@
Thus, our definition includes three quantitative criteria ($P<30$ ms, $P_{\rm b}\lesssim 10$ d and $M_{\rm c,min}>0.004$ M) and one qualitative criterion assessing the nature of the companion star (non-degenerate or semi-degenerate). The latter is necessary to exclude MSP-WD systems with fully degenerate companions, which can overlap with spiders in $P$, $M_{\rm c,min}$ and sometimes $P_{\rm b}$. As the sample grows – and as discussed below – peculiar systems are being discovered. Examples include the RB PSR J1932+2121 with $P=14.2$ ms (D. Misra et al., 2025b), the RB-hu PSR J1947-1120 with $P_{\rm b}=10.3$ d (J. Strader et al., 2025), and the RB PSR J1908+2105 with $M_{\rm c,min}=0.05$ M (J. A. Simpson et al., 2025).
@@ 2 Spider-Cat: a catalog of compact binary millisecond pulsars
We present the spider catalog (SpiderCat), which compiles spin, orbital, and multi-wavelength information for all publicly known spider systems in the Galactic field, excluding those in globular clusters [^1] For globular cluster MSPs, see https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html.. The paper is accompanied by an online database [^2] https://astro.phys.ntnu.no/SpiderCAT, where the information presented here is available for easy access.
In addition to confirmed spider systems, we also include spider candidates (denoted RBc and BWc), which exhibit similar characteristics to confirmed systems but lack definitive pulsation detections. These candidates are identified based on several observational signatures, as discussed in Section 1.1. Although pulsations have not yet been detected in these systems, their multi-wavelength properties strongly suggest a classification as spider MSPs  [^3] Indeed, follow-up observations of several spider candidates have detected pulsations, e.g., in the BW PSR J1311-3430 (R. W. Romani, 2012; H. J. Pletsch et al., 2012), the RB PSR J2339-0533 (R. W. Romani & M. S. Shaw, 2011a; H. J. Pletsch & C. J. Clark, 2015) or the RB PSR J0212+5321 (M. Linares et al., 2017; K. I. Perez et al., 2023). .
In the following sections, we describe the source sample (Sec. 2.1) and catalogue parameters (Sec. 2.2). We present the Galactic distribution of spiders in Section 3.1, the spider $P-\dot{P}$-diagram in Section 3.2, and the $M_{\rm c,min}-P_{\rm b}$ diagram in Section 3.3. In Section 3.4, we discuss and tabulate distance estimates for spider systems. Section 3.5 presents the multiwavelength counterparts ($\gamma$-ray/X-ray/optical/IR), including their fluxes, estimated luminosities, and apparent and absolute magnitudes. We discuss potential spin-orbit grouping of spiders in Section 3.6 and conclude in Section 4.

### v4

@@ Abstract
We present SpiderCat, a multiwavelength catalog of all publicly known compact binary millisecond pulsars (MSPs) in the Galactic field. These systems, colloquially known as “spiders,” consist of neutron stars in tight orbits with low-mass companions, which are gradually ablated by the pulsar wind. SpiderCat includes both primary subclasses—redbacks and black widows—distinguished by companion mass, as well as candidates and peculiar systems such as transitional, huntsman, and tidarren MSPs. As of this initial release, SpiderCat contains 111 entries: 30 redbacks, 50 black widows, two huntsmans, 23 redback candidates, five black widow candidates, and one huntsman candidate. In this paper, we compile and summarize key parameters for each system, including spin and orbital properties, and multiwavelength data from radio, optical, X-ray, and $\gamma$-ray observations. An interactive, publicly accessible web interface, at https://astro.phys.ntnu.no/SpiderCAT, enables exploration and visualization of the data. The rapid growth of the number of known spiders, accelerated by the Fermi Large Area Telescope survey and its ability to identify MSPs in $\gamma$-rays, has opened the door to population-level studies. Utilizing SpiderCat, we analyze trends in spin period, orbital period, companion mass, emission properties, and spatial distribution. SpiderCat serves as a dynamic, multiwavelength repository for this unique class of binary pulsars, facilitating new discoveries and constraints on pulsar evolution, particle acceleration, and the neutron star equation of state.
[…]
@@ 1 Introduction
Before the Fermi $\gamma$-ray Space Telescope was launched in 2008, only seven $\gamma$-ray pulsars had been confidently detected (by the Energetic $\gamma$-Ray Experiment Telescope (EGRET), between 1991 and 2000), all of which were young and slowly rotating (spin period $P>30$ ms; D. J. Thompson, 2008). The Large Area Telescope (LAT) onboard Fermi, with an effective area nearly 10 times larger than EGRET’s, has proven to be an efficient pulsar-discovery machine (W. B. Atwood et al., 2009). Based on 12 yr of all-sky survey LAT data, Fermi’s Third Pulsar Catalog (3PC) includes 294 $\gamma$-ray pulsars (and 45 candidates; D. A. Smith et al., 2023). Roughly half of these 3PC $\gamma$-ray pulsars are millisecond pulsars (MSPs): “old,” rapidly spinning ($P<30$ ms) neutron stars in binary systems, with relatively low magnetic fields ($B\sim 10^{8}-10^{9}$ G), that have been “recycled” through the accretion of matter (M. A. Alpar et al., 1982). Because MSPs have a higher Galactic scale height than young pulsars (D. R. Lorimer, 2013), they are more easily discovered in LAT surveys far from the Galactic plane, where $\gamma$-ray diffuse emission does not obscure faint point sources.
@@
Many of the new Fermi-LAT pulsar discoveries have been identified as compact binary MSPs, characterized by short orbital periods (typically $P_{\rm b}\lesssim 1$ d) and nondegenerate or semidegenerate companion stars (see definitions in Section 1.2). In such compact systems, interactions between the pulsar and its companion can be extreme; the outer layers of the companion star may be ablated leading to occultations or “eclipses” of the radio pulsar (A. S. Fruchter et al., 1988). These destructive effects have inspired the cannibalistic spider nicknames for compact binary MSPs—such as black widows (BWs) and redbacks (RBs)—commonly referred to in the literature (and hereinafter) as spider pulsars, spider MSPs, or simply “spiders” (e.g., M. S. E. Roberts, 2011).
@@
Pulsars represent the most numerous Galactic class of $\gamma$-ray sources. They are characterized by low $\gamma$-ray variability and high spectral curvature, which sets them apart from blazars (the most numerous extragalactic class) that present typically higher variability and power law spectra. These characteristics enable an efficient selection of pulsar candidates among thousands of unidentified Fermi-LAT sources (>>7000 cataloged sources, with >>2000 still unidentified; S. Abdollahi et al., 2022). Thus, the Fermi-LAT point source catalogs have been a treasure trove for spider discoveries over the past 15 yr (Section 1.1).
@@
An “elite” subset of compact binary MSPs (belonging to the RB subtype, see below) have been observed in the so-called disk state, where they form an accretion disk and the radio pulsar becomes undetectable (Z. Wang et al., 2009; D. de Martino et al., 2015). These so-called transitional MSPs (tMSPs; labeled “tr” below) show optical disk emission lines and a characteristic X-ray variability pattern known as X-ray mode switching (M. Linares et al., 2014). In the pulsar state, when the radio pulsar is detected, spiders show lower luminosities and orbital modulation (with different origins) throughout the electromagnetic spectrum, from radio to $\gamma$-rays. These properties are used to identify new spiders, as described in Section 1.1.
@@
As a consequence of the compact orbits, an intrabinary shock (IBS) is thought to form between the pulsar and companion winds in spider MSPs (J. Arons & M. Tavani, 1993; E. S. Phinney et al., 1988). This shock produces non-thermal (synchrotron) X-ray emission that is modulated with the orbit (e.g., B. W. Stappers et al., 2003; R. H. H. Huang et al., 2012; R. W. Romani & N. Sanchez, 2016; Z. Wadiasingh et al., 2017; C. J. T. van der Merwe et al., 2020). The IBS can also be an efficient site for particle acceleration (e.g., A. K. Harding & T. K. Gaisser, 1990; L. Sironi & A. Spitkovsky, 2011; J. Cortés & L. Sironi, 2022, 2024, 2025), potentially contributing to the cosmic-ray positrons detected on Earth (M. Linares & M. Kachelrieß, 2021). During the gigayear-long recycling process, neutron stars in low-mass X-ray binaries can accrete up to about 1 M (e.g., D. Misra et al., 2025b). As a result, MSPs in general—and spiders in particular—are thought to harbor massive neutron stars. Indeed, several studies have found evidence for massive neutron stars from dynamical measurements of compact binary MSPs (e.g., M. Linares et al., 2018; J. Strader et al., 2019; R. W. Romani et al., 2022).
@@
For the reasons above, the field of compact binary MSPs has experienced a boom in the last decade. We now know of more than 100 spiders, enabling meaningful population studies. While existing pulsar catalogs (Section 2.1) focus on radio timing properties, there is no public, up-to-date repository of multiwavelength information for spider MSPs (to the best of our knowledge). Thus, the time is ripe to collect and analyze the spin, orbit and multiwavelength properties of all known Galactic-field spiders. This is the goal of the present work.
@@ 1.1 Spider Discoveries
Figure 1: Histogram showing the discovery year of spider systems included in SpiderCat, color coded by spider subclass. The discovery year is defined as the year in which a paper was published confirming the system as a spider or spider candidate, or reporting the detection of millisecond pulsations from the source. The first four spider systems are highlighted. The dramatic increase in discoveries following the launch of the Fermi $\gamma$-ray Space Telescope reflects its critical role in guiding targeted radio (e.g., 350 MHz Green Bank Telescope (GBT) and 327 MHz Arecibo surveys) and multiwavelength surveys.
@@
Thanks to a flurry of Fermi-driven discoveries, the spider population has grown from a rara avis (with only four systems known before 2008; see Figure 1) into an important subclass of MSPs (109 systems presented here as of 2025). Indeed, spiders now constitute 15%–20% of the 552 MSPs currently known in the Galactic disk (E. Ferrara & D. Lorimer, 2025, where the higher end of the range includes candidate systems). The methods that have been used to discover new spiders include the following.
  1. 1.
@@
    Wide-field radio timing surveys (e.g., the Galactic Plane Pulsar Snapshot (GPPS) survey with the Five-hundred-meter Aperture Spherical Telescope (FAST); P. F. Wang et al. 2025; see also A. J. Faulkner et al. 2004). These surveys typically observe each sky location once or only a few times, with relatively short integrations. Because in spider systems the pulsar is occulted during a large fraction of the orbit (20%–60%), these surveys are suboptimal for discovering them. Follow-up observations are required to measure orbital parameters.
  2. 2.
@@
    $\gamma$-ray-guided radio timing searches (e.g., by the Pulsar Search Consortium using the GBT; J. W. T. Hessels et al. 2011; P. S. Ray et al. 2012; see also E. D. Barr et al. 2013). This method targets specific unidentified $\gamma$-ray sources using longer and/or more numerous radio pointings, after selecting pulsar candidates based on their $\gamma$-ray properties (as discussed above). These searches have been particularly effective in discovering new spiders, which are readily identified by their radio “eclipses” or occultations.
  3. 3.
@@
    $\gamma$-ray-guided optical searches (e.g., the Compact Binary Pulsar Search; M. Turchetta et al. 2024; see also R. W. Romani & M. S. Shaw 2011; A. K. H. Kong et al. 2012; M. Linares et al. 2017; C. Lu et al. 2025). Compared to the more common binary MSPs with white dwarf (WD) companions, spider MSPs with main-sequence and brown dwarf companions are typically more luminous in the optical–IR band. Their compact orbits result in large-amplitude orbital modulations due to the gravitational distortion and irradiation of the companion star, as viewed at different orbital phases. These features enable optical–IR searches for spider candidates, especially the optically brighter RB subclass (Sections 1.2 and 2.1), by targeting their characteristic flux (and radial velocity) variations. The candidate sources are typically selected based on $\gamma$-ray properties (low variability and highly curved spectra).
  4. 4.
@@
    Blind $\gamma$-ray searches (e.g., H. J. Pletsch et al. 2012; see also L. Nieder et al. 2020). The Fermi-LAT has been surveying the full sky since 2008, providing a rich dataset for blind pulsar searches. These searches fold $\gamma$-ray photons over a range of spin and orbital parameters to detect pulsations from a given source or sky location. Although computationally intensive, especially for MSPs with unknown orbital parameters, they can be aided by prior knowledge of such parameters, e.g., the orbital period (C. J. Clark et al., 2021). In addition, some spiders have shown orbital modulation (H. An et al., 2020) and eclipses (C. J. Clark et al., 2023a) in their $\gamma$-ray light curves.
  5. 5.
@@
    X-ray searches (e.g., S. Bogdanov & J. P. Halpern 2015; see also F. Coti Zelati et al. 2019). X-ray properties can also help identify new spider systems. In particular, X-ray mode switching has been used to discover new tMSP candidates in the disk state. In the pulsar state, features such as the X-ray luminosity ($\sim 10^{30}-10^{33}$ erg s-1), spectral shape (photon index $\simeq 1.5-3$), or orbital modulation from the IBS (often double peaked with a flux maximum near pulsar inferior conjunction) can aid in identifying spider systems.
@@ 1.2 Spider Definition and Taxonomy
  1. 1.
    a pulsar spin period $P<30$ ms (the majority of identified systems have spins $1.5<P<5.0$ ms); [^1] This limit is a convention commonly used in the literature for MSPs (e.g., D. Bhattacharya & E. P. J. van den Heuvel 1991; D. R. Lorimer 2008) and corresponds to a sparsely populated region in the spin period distribution separating the main pulsar and MSP populations (see Section 3.2).
  2. 2.
    an orbital period $P_{\rm b}\lesssim 10$ days (most systems have $P_{\rm b}\lesssim 1$ day, indicative of compact orbits; the exceptions are the so-called huntsman (HM) pulsars, defined below, which have wider orbits of 2–10 days); and
  3. 3.
    a nondegenerate or semidegenerate companion star, inferred from (i) the presence of radio eclipses, (ii) the detection of X-ray mode switching (in tMSPs), (iii) the optical–IR spectra of the companion or accretion disk, and (iv) the detection and properties of X-ray and/or optical orbital modulation (see details above).
Spiders are divided into two broad subclasses, based on the minimum companion mass ($M_{\rm c,min}$):
  1. 1.
@@
For completeness, we also include a related class of systems known in the literature as HMs:
  1. 1.
    HM pulsars are systems with giant companions and $P_{\rm b}$ in the 2–10 day range (J. Strader et al., 2015; S. J. Swihart et al., 2017; J. Strader et al., 2025). Due to their luminous companions and wide orbits, the companion stars are very weakly irradiated and are likely not ablated. However, these companions have masses similar to those found in RB systems and exhibit evidence of IBSs, akin to those seen in canonical spider systems (S. J. Swihart et al., 2018).
Thus, our definition includes three quantitative criteria ($P<30$ ms, $P_{\rm b}\lesssim 10$ days and $M_{\rm c,min}>0.004$ M) and one qualitative criterion assessing the nature of the companion star (nondegenerate or semidegenerate). The latter is necessary to exclude MSP–WD systems with fully degenerate companions, which can overlap with spiders in $P$, $M_{\rm c,min}$, and sometimes $P_{\rm b}$. As the sample grows—and as discussed below—peculiar systems are being discovered. Examples include the RB PSR J1932+2121 with $P=14.2$ ms (D. Misra et al., 2025a) and the RB PSR J1908+2105 with $M_{\rm c,min}=0.05$ M (J. A. Simpson et al., 2025).
@@ 2 Spider-Cat: A Catalog of Compact Binary Millisecond Pulsars
We present the spider catalog (SpiderCat), which compiles spin, orbital, and multiwavelength information for all publicly known spider systems in the Galactic field, excluding those in globular clusters. [^2] For globular cluster MSPs, see https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. The paper is accompanied by an online database, [^3] https://astro.phys.ntnu.no/SpiderCAT where the information presented here is available for easy access.
In addition to confirmed spider systems, we also include spider candidates (denoted RB candidates (RBcs) and BW candidates (BWcs)), which exhibit similar characteristics to confirmed systems but lack definitive pulsation detections. These candidates are identified based on several observational signatures, as discussed in Section 1.1. Although pulsations have not yet been detected in these systems, their multiwavelength properties strongly suggest a classification as spider MSPs. [^4] Indeed, follow-up observations of several spider candidates have detected pulsations, e.g., in the BW PSR J1311-3430 (R. W. Romani, 2012; H. J. Pletsch et al., 2012), the RB PSR J2339-0533 (R. W. Romani & M. S. Shaw, 2011; H. J. Pletsch & C. J. Clark, 2015) or the RB PSR J0212+5321 (M. Linares et al., 2017; K. I. Perez et al., 2023).
In the following sections, we describe the source sample (Section 2.1) and catalog parameters (Section 2.2). We present the Galactic distribution of spiders in Section 3.1, the spider $P-\dot{P}$-diagram in Section 3.2, and the $M_{\rm c,min}-P_{\rm b}$ diagram in Section 3.3. In Section 3.4, we discuss and tabulate distance estimates for spider systems. Section 3.5 presents the multiwavelength counterparts ($\gamma$-ray, X-ray, optical, and IR), including their fluxes, estimated luminosities, and apparent and absolute magnitudes. We discuss potential spin-orbit grouping of spiders in Section 3.6 and conclude in Section 4.